OpenStudy (winsteria):

FAN AND MEDAL Use slope-intercept form, y = mx + b to find the equation of the line that passes through the points (−6, 1) and (3, 4) 1)y = –3x + 5 2)y = 3x – 5 3)y=1/3x+3 4)y=-1/3x+5

OpenStudy (winsteria):

@bibby

OpenStudy (bibby):

find the slope first

OpenStudy (winsteria):

so...

OpenStudy (winsteria):

but that's none of the choices

OpenStudy (bibby):

slope of points between (x1,y1) and (x2,y2) \(\dfrac{y_2-y_1}{x_2-x_1}=m\) \(m=\dfrac{4-1}{3-(-6)}=\dfrac{3}{9}=\dfrac{1}{3}\) \(y-y_1=m(x-x_1)\\ y-4=\dfrac{1}{3}(x-3) \\y-4=\dfrac{1}{3}x-1 \\y=\dfrac{1}{3}x+3\)

OpenStudy (winsteria):

thank you